# Effective Online Exam Proctoring by Combining Lightweight Face Detection and Deep Recognition

### v2

@@ VI-C A Case Study of iExam’s Post-exam Analysis Report
Temporal Presence and Absence Analysis. In addition to overall recognition frequency, the system generates fine-grained temporal plots, such as those in Fig. 12, which display the second-by-second presence status for each participant throughout the exam. Each bar segment corresponds to one second where a student was present or absent, allowing instructors to pinpoint the exact times when a student disappeared from view. This feature is particularly useful for cross-checking suspected incidents of absence or face swapping, providing an objective and auditable record for subsequent review.
Figure 12: Summary of recognition distribution. The visualization presents a timeline showing each student’s presence or absence at every second throughout the exam period. To safeguard privacy, real names were anonymized using generic labels.
Consecutive Absence Detection. Finally, as summarized in Fig. 13, the system aggregates the total duration of consecutive absences for each student, highlighting cases where a participant was missing for an extended period. Longer blocks of continuous absence are considered more suspicious and may warrant targeted investigation or follow-up with the corresponding student. This information is presented both as a summary bar chart and can be cross-referenced with the original video for validation.
Figure 13: Consecutive disappear time for testing video. Each bar indicates the longest continuous absence interval detected for each student. For privacy protection, actual names were substituted with generic labels.
@@
By providing automated, interpretable, and time-aligned visual feedback, iExam’s post-exam analysis report enables instructors to efficiently verify student presence, investigate anomalies, and maintain exam integrity with minimal manual effort. The reporting system is scalable to longer videos and larger cohorts, and can be customized or extended for more detailed event logging or alerting as needed.
VII Related Work
Traditional Face Recognition Approaches. Early face recognition research primarily focused on holistic and feature-based statistical methods. The Eigenface approach, based on Principal Component Analysis (PCA) [turk1991eigenfaces, ramadhani2017human], projects face images onto a lower-dimensional subspace spanned by principal components, capturing the main variation across training images. Fisherface [chan2008face] applies Linear Discriminant Analysis (LDA) to maximize between-class separability while minimizing within-class variance, providing robustness against lighting and expression changes. Another commonly used traditional method, Local Binary Patterns Histograms (LBPH) [LBPH], encodes local texture information by thresholding pixel neighborhoods, then aggregates histograms for recognition. Despite their historical significance and efficiency, these methods suffer from degraded performance under varying lighting, pose, and occlusion conditions, and lack the capacity to model non-linear correlations in high-dimensional face data.
Haar-like Features and Classical Detection. Viola and Jones [viola2001rapid] introduced a cascade classifier based on Haar-like features and Adaboost, which became a widely used standard for real-time face detection. Haar-like features capture simple intensity differences between adjacent rectangular regions (e.g., eyes vs. cheeks), leveraging the biological symmetry of faces for efficient object localization. Such approaches, while fast, are limited in their expressiveness and tend to be brittle in unconstrained environments [chen2015enhancing].
OpenCV Face Recognizers. The OpenCV library provides classical, well-established algorithms for face recognition, including the Eigenface, Fisherface, and LBPH methods [Eigenface, Fisherface, LBPH]. These approaches reply primarily on traditional feature extraction and dimensionality reduction techniques rather than deep neural networks. While these implementations are efficient and computationally lightweight, which makes them suitable for small-scale or embedded applications, their performance degrades in complex, real-world scenarios. Online proctoring scenarios introduce challenges such as large pose variations, inconsistent illumination, motion blur, partial occlusion, and low-resolution webcam feeds, all of which severely affect the discriminative power of classical methods. Moreover, these algorithms depend heavily on the quality of face alignment and are limited in their ability to generalize across individuals or sessions without extensive manual calibration. These limitations have motivated the shift toward deep learning-based face recognition models, which often require more computational resources but deliver significantly improved accuracy and robustness. To address these trade-offs, our system incorporates automated model selection, enabling dynamic switching between the classical algorithm, i.e., Haar-cascade method, and Deep Convolutional Neural Networks (DCNNs) based on runtime resource availability and performance requirements.
Deep Learning and CNN-based Approaches. Recent advances in deep learning have revolutionized the field of face recognition and detection. Deep Convolutional Neural Networks (DCNNs) [lecun1998gradient, krizhevsky2012imagenet, he2016deep] learn hierarchical feature representations directly from raw images, automatically capturing complex, non-linear facial characteristics such as pose, illumination, and expression variations. Landmark architectures such as AlexNet [krizhevsky2012imagenet], GoogleNet [szegedy2015going], and ResNet [he2016deep] have demonstrated that increasing network depth and introducing residual connections significantly improve generalization and robustness across large-scale visual recognition benchmarks. Building on these foundations, specialized face recognition models, such as FaceNet [schroff2015facenet] and ArcFace [deng2019arcface], achieve state-of-the-art accuracy by learning discriminative embeddings that map facial images into compact feature spaces. In the context of online proctoring, DCNN-based models provide the accuracy, adaptability, and robustness needed to handle diverse webcam feeds and dynamic exam environments.
Face Detection Algorithms. Several modern deep learning-based detection algorithms have been developed, including Multi-task Cascaded Convolutional Networks (MTCNNs) [zhang2016joint], Single Shot-Multibox Detector (SSD) [liu2016ssd], and DCNN models integrated with frameworks such as OpenCV. These algorithms have proven highly effective in challenging real-world conditions and they can accurately localize faces in complex scenes and are less sensitive to pose and illumination changes compared to classical methods. Our system evaluates and leverages the performance and efficiency trade-offs of these mainstream detectors for both real-time and post-exam analysis.
Optical Character Recognition (OCR) in Human Identity Verification. OCR plays a pivotal role in scenarios where automatic identity extraction from video is needed. State-of-the-art OCR engines, such as Google’s Tesseract (used via PyTesseract in Python), perform multi-stage processing including image acquisition, denoising, binarization, segmentation, feature extraction, and text recognition. They are widely adopted in academic and industrial applications for extracting textual identity cues from screen overlays or exam platforms [tesseract_github]. Recent research also explores integrating OCR with face recognition to enable automatic ground-truth labeling in large-scale datasets [arlazarov2018midv, bulatov2021midv], a technique we adopt to facilitate hands-free, dynamic identity annotation in online exams.
@@ VII Related Work
Automated Proctoring Systems. A number of recent works and commercial solutions (e.g., ProctorU and open-source frameworks) aim to automate the monitoring of remote exams via video analysis, using face detection, face recognition, gaze estimation, and anomaly detection [proctoru_official_website, butler2020systematic]. However, few systems provide fine-grained, two-stage separation of lightweight real-time detection and deep post-exam analysis as in iExam. Our method uniquely addresses the trade-off between speed, privacy, and post-exam integrity verification by combining client-side lightweight detection with deep learning-powered offline recognition and reporting.
In summary, while classical statistical and feature-based methods laid the groundwork for face analysis, the state-of-the-art has shifted toward deep learning-based models for robust performance under realistic, unconstrained conditions. Our approach leverages this paradigm shift, integrating efficient DCNN-based detection and recognition, automated OCR labeling, and tailored analysis for online exam proctoring, to deliver a not only accurate and scalable but also privacy-preserving solution.
VIII Conclusion
@@
This paper presented iExam, an effective online exam proctoring system that combines lightweight real-time face detection with deep learning-based post-exam face recognition. By decoupling instant monitoring from in-depth analysis, iExam achieves both efficiency and high accuracy while protecting user privacy. Experimental results demonstrate that the system delivers robust performance on real-world exam videos, enabling reliable detection of student presence and identification of suspicious behaviors. Our findings suggest that iExam offers a practical and scalable solution for secure and transparent online exam supervision.

### v3

@@ IV-C A Case Study of iExam’s Post-exam Analysis Report
Temporal Presence and Absence Analysis. The system generates second-by-second presence plots for each participant, as in Fig. 7(b). Each segment represents one second of presence or absence, enabling instructors to pinpoint exact times a student disappeared from view. This facilitates cross-checking suspected incidents such as absence or face swapping and provides an auditable record for review.
Consecutive Absence Detection. As in Fig. 7(c), iExam aggregates consecutive absence durations per student, highlighting extended absences that may require follow-up. Summaries are presented as bar charts and can be cross-referenced with video recordings for validation.
@@
By delivering automated, interpretable, and time-aligned visual feedback, iExam’s post-exam reports allow instructors to efficiently verify presence, investigate anomalies, and maintain exam integrity, while scaling to longer exams and larger cohorts.
V Related Work
Deep Learning-based Face Detection and Recognition. Recent advances in deep learning have substantially improved performance. Convolutional Neural Networks (CNNs) [howard2017mobilenets, huang2017densely, SqueezeNet, szegedy2015going] and Deep CNNs (DCNNs) [lecun1998gradient, krizhevsky2012imagenet, he2016deep] achieve high accuracy and robustness, particularly when pre-trained on large-scale face datasets. Several models [zhang2016joint, wang2023efficientface] demonstrate low detection error rates and fast inference, making them suitable for automated proctoring scenarios.
@@ V Related Work
Automated Proctoring Systems. Recent academic and commercial systems leverage video-based analysis for remote exam monitoring, incorporating face detection, recognition, and anomaly detection [proctoru_official_website, butler2020systematic]. However, most approaches lack a clear separation between real-time monitoring and post-exam analysis. In contrast, iExam adopts a two-stage design that combines lightweight client-side detection with deep offline recognition, explicitly balancing efficiency, privacy, and exam integrity verification.
VI Conclusion
@@
This paper presented iExam, an effective online exam proctoring system that combines lightweight real-time face detection with deep learning-based post-exam face recognition. By decoupling instant monitoring from in-depth analysis, iExam achieved both efficiency and high accuracy while protecting user privacy. Experimental results demonstrated that the system delivers robust performance on real-world exam videos, enabling reliable detection of student presence and identification of suspicious behaviors. Our findings suggested that iExam offers a practical and scalable solution for secure and transparent online exam supervision.